Freon in air conditioning unit has a volume of 0.30L. It is allowed to function in a room, where the pressure is about 628 mm of Hg, assuming that the temperature is in a constant state. Find the final pressure of Freon when its volume is increased to 0.95L.


1
Expert's answer
2022-03-03T07:19:01-0500

Boyle's law

P1V1 = P2V2

V1 = 0.30L, V2 =0.95L

P1 = 628 mm of Hg

P2 = P1V1/V2 = 628 * 0.3/0.95 = 198 mm of Hg

Answer: P2 = 198 mm of Hg